OpenStudy (anonymous):

Evaluate (a^2b - 3ab) / b

OpenStudy (chaise):

(a^2b - 3ab) / b = b(a^2x-3a)/b = a^2x-3a = a(ax-3)

OpenStudy (chaise):

Disregard the x's there >.<

OpenStudy (anonymous):

but no x was given

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

why hv u changed bs position